OpenStudy (anonymous):

Solve the system of equations. 2x - 4y = 3 -3x + 5y = 1

OpenStudy (dls):

Multiply 1st equation by 3 and 2nd by 2 nd solve simultaneously

OpenStudy (anonymous):

would i multiply 2x by three , 4y by 3 and 3 by 3? or how would i multiply it

OpenStudy (dls):

every term

OpenStudy (anonymous):

so then it'd be 6y - 12y = 6?

OpenStudy (dls):

2nd one too

OpenStudy (anonymous):

multiply the second one by 2?

OpenStudy (anonymous):

or 3 also

OpenStudy (dls):

Sorry for late replies,kinda busy..let me do it for u now

OpenStudy (dls):

See basically,u have to find a common term in both so we can eliminate one term..so that we have a single equation variable to solve :)

OpenStudy (dls):

If you multiply \[2x - 4y = 3 \] by 3 and \[-3x + 5y = 1\] by 2 u willget 6x in both so when u solve both of them simultaneously..  solve for y,substitute in any eq. and find x.